OpenStudy (anonymous):

A solid weighs 277.5 gm in air and 212.5 gm when totally immersed in the liquid of density 0.9 gm/cm^3. Calculate the density of the solid. Ans:2.05 gm/cm^3 plzzz help meteach the process to solve dis..plzz

OpenStudy (anonymous):

Do u know the formula for Density??

OpenStudy (anonymous):

ya..it mass per volume!

OpenStudy (anonymous):

Then find the volume occupied by solid in that liquid using the formula!!

OpenStudy (souvik):



OpenStudy (anonymous):

bt hw?

OpenStudy (anonymous):

Then find the density of solid by taking its mass as 227.5 gm. ....as volume of a solid doesnt change(in this case)

OpenStudy (anonymous):

Use d=m/v;where m=212.5 gm and d=0.9 gm/cm^3.Get v!1

OpenStudy (anonymous):

Substitute this v in d(solid)=mass(solid)/(v)volume(solid)

OpenStudy (anonymous):

Gtin it???

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

So can u cntnu??

OpenStudy (anonymous):

yup tq :D

OpenStudy (anonymous):

Ur Welcome.Cheer!!:)

OpenStudy (souvik):

let V be the volume of the solid and \(\rho\) be the density of the solid so buoyant force=weight in air-weight in water =277.5-212.5=65 gm-wt now the buoyant force = V*density of liquid find V from here... now the density of the solid=m/V

OpenStudy (anonymous):

quite late bt tq 4 ur help souvik.. :)

OpenStudy (souvik):

welcome...:)

OpenStudy (anonymous):

Is my method wrong souvik??Pls do tell.I dont want sum1 to learn a wrong method!1

OpenStudy (souvik):

the way u find the volume of the solid....dividing the weight by the density of liquid ....i think its wrong..!

OpenStudy (anonymous):

@sweta.aryal

OpenStudy (anonymous):

Which method gives u the answer??Pls check

OpenStudy (anonymous):

@souvik I think neither of our methods give her the answer!!:(

OpenStudy (anonymous):

Mine comes to around 1.19,while urs to 3.86

OpenStudy (souvik):

only @sweta.aryal can tell if she wrote the correct answer or not..!!!

OpenStudy (anonymous):

True